OpenStudy (gmail.comlover):

the vertices of the triangle are (-3,1) (1,1) and (1, -2) (coordination graph) What is the perimeter of the triangle?

OpenStudy (nincompoop):

determine the length from one point to another

OpenStudy (nincompoop):

assume the distance between two points A and B, \(distance = d = \sqrt{A^2+B^2} = \large \sqrt{(x_f-x_i)^2+(y_f-y_i)^2} \)

OpenStudy (gmail.comlover):

okay thank you so much! that was helpful